$13,138 is invested, part at 9% and the rest at 7%. If the interest earned from the amount invested at 9% exceeds the interest earned from the amount invested at 7% by $537.46, how much is invested at each rate? (Round to two decimal places if necessary.)

Answer by jorel555(1290)

Let n be the amount invested at 9%. Then:
.09n-537.46=.07(13138-n)
9n-53746=91966-7n
16n=145712
n=9107
$9107 was invested at 9%; and $4031 was invested at 7%. ☺☺☺☺